Recent Log Entries

70 uA from Hall A to FC1 with FLIP LEFT

Log Number: 3684470Logbook(s): Entry Makers: suleimanBody: 

Transmission is about 85% with Gun Supply current (82 uA) and 92% with PCup current (76 uA). Relative orbit looked fine and tried to adjust A laser phase but with minimal effect (extra 1-2 uAmps).

iocha48 restored

Log Number: 3684467Logbook(s): Tags: Body: restorelib 2.0.1 Target restore date was 2019-04-04 19:22:23 The output from the myRestore execution was as follows: iocha48 magnet_MAG@2019-04-04 19:22 12(12) restored Pause: 2 second(s) magnet_HVAL@2019-04-04 19:22 3(3) restored Pause: 2 second(s) magnet_LOOP@2019-04-04 19:22 3(3) restored Pause: 2 second(s) magnet_CMPQ@2019-04-04 19:22 6(6) restored system@2019-04-04 19:22 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocha48 (success) Procedure return status: Success

UITF flow switch FC4/A5/A6 ready

Log Number: 3684464Logbook(s): Body: 

UITF flow switch for FC4/A5/A6 was tested and adjusted this morning. It is ready for beam. Here is a list of all the current tested and adjusted flow switches in UITF right now.

FC1/MS flow .40gpm trip set <.38gpm

A3/A4 flow .80gpm trip set <.50gpm

FC2/D1 flow .98gpm trip set <.50gpm

FC3 flow 1.40 gpm trip set <.70 gpm

FC4/A5/A6 flow .62 gpm trip set <.50 gpm

D2/D3 not testing/adjusted yet but is in progress today.

iocitf1 restored

Log Number: 3684459Logbook(s): Tags: Body: restorelib 2.0.1 Target restore date was 2019-05-31 07:02:49 The output from the myRestore execution was as follows: iocitf1 system@2019-05-31 07:02 2(2) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ued -A -s -i iocitf1 (success) Procedure return status: Success

COMMENT ON Remove Hall A Dump Aperture Plate

Log Number: 3684456Logbook(s): Body: 

New metal seal due to arrive today, installation will happen
as soon as we have it in hand today, the plan would then be
to pump down over the weekend and install the shroud on
Monday if everything looks good.

IPM1F01 is broken

Log Number: 3684455Logbook(s): Body: 

Checking LERF 4ch and SEE bpms this morning and find that IPM1F01 4ch bpm is broken. I don't know what happened to it but this bpm is failing local X/Y alpha's and looks terrible with Supercal. Definitely needs to be fixed.

Taking the NL to Controlled Access

Log Number: 3684454Logbook(s): Entry Makers: aikenBody: 

FC2 has remained in (IFY0L03T), and no gradient in the INJ/NL RF except for the quarter cryo (*GMES).

Pages